OpenStudy (anonymous):

Im stuck on this problem , Help please ? Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3).

OpenStudy (anonymous):

When two lines are parallel their slopes are the same. Use this formula, \[(y - y_1) = m(x-x_1)\] Then solve for y

OpenStudy (anonymous):

I just used that formula and Im still not understanding.

OpenStudy (anonymous):

What are you confused about?

OpenStudy (anonymous):

Parallel lines have same slope then the required equation becomes y=7x+k k can be evaluated from the point given put x and y here

OpenStudy (anonymous):

How do I incorporate that into this equation? I plugged everything into the formula (y-y1)=m(x-x1) and got y+3=7(x-1)

OpenStudy (anonymous):

Now what do I do? :/

OpenStudy (anonymous):

-3=7(1)+k implies k=-10 Required equation is y=7x-10 where slope is 7 and intercept is -10

OpenStudy (anonymous):

Atchyut please help me with the question we started before please

OpenStudy (anonymous):

Can u direct me to that question????

OpenStudy (anonymous):

Now we have to use algebra to manipulate the equation into slope-intercept form. \[y +3 = 7(x-1)\]Now, let's subtract 3 from both sides.\[y = 7(x-1) - 3\]Now, let's distribute the 7. \[y = 7x - 7 - 3\]Now, let's combine like terms. \[y = 7x - 10\]

OpenStudy (anonymous):

Oh that's simple , thankyou :) !